# Fourier Circuits in Neural Networks: Unlocking the Potential of Large Language Models in Mathematical Reasoning and Modular Arithmetic

### v1

@@ 3.3 Preliminary
Condition 3.10 (Condition C.1 in page 8 in [55]).
We have $g^{\prime}(\theta^{*},x,y)=g(\theta^{*},x,y)$ for all $(x,y)\in\operatorname{spt}(q^{*})$. It means:
$\{y^{\prime}\in\mathcal{Y}\backslash\{y\}:\tau(x,y)[y^{\prime}]>0\}\subseteq\underset{y^{\prime}\in\mathcal{Y}\backslash\{y\}}{\arg\max}f(\theta^{*},x)[y^{\prime}].$
Hence, by satisfying these conditions, we will concentrate on describing the class-weighted maximum-margin solutions.
@@ 4 Main Result
We characterize the Fourier features to perform modular addition with $k$ input in the one-hidden-layer neuron network. We show that every neuron only focus on a distinct Fourier frequency. Additionally, within the network, there is at least one neuron of each frequency. When we consider the uniform class weighting so that $\mathcal{L}_{\lambda}(\theta)$ is based on $\tau$,
$\tau(a_{1},\dots,a_{k})[c^{\prime}]:=1/(p-1)~{}~{}\forall c^{\prime}\neq a_{1}+\dots+a_{k},$ (6)
we have the following main result:
Theorem 4.1 (Informal version of Theorem F.2).
Let $f(\theta,x)$ be the one-hidden layer networks defined in Section 3. If $m\geq 2^{2k-1}\cdot\frac{p-1}{2}$, then the max $L_{2,k+1}$-margin network satisfies:
@@
    $u_{1}(a_{1})$ $=\beta\cdot\cos(\theta_{u_{1}}^{*}+2\pi\zeta a_{1}/p)$
    $u_{2}(a_{2})$ $=\beta\cdot\cos(\theta_{u_{2}}^{*}+2\pi\zeta a_{2}/p)$
    $\dots$
    $u_{k}(a_{k})$ $=\beta\cdot\cos(\theta_{u_{k}}^{*}+2\pi\zeta a_{k}/p)$
    $w(c)$ $=\beta\cdot\cos(\theta_{w}^{*}+2\pi\zeta c/p),$
@@
Figure 1: Cosine shape of the trained embeddings (hidden layer weights) and corresponding power of Fourier spectrum. The two-layer network with $m=2944$ neurons is trained on $k=4$-sum mod-$p=47$ addition dataset. We even split the whole datasets ($p^{k}=47^{4}$ data points) into the training and test datasets. Every row represents a random neuron from the network. The left figure shows the final trained embeddings, with red dots indicating the true weight values, and the pale blue interpolation is achieved by identifying the function that shares the same Fourier spectrum. The right figure shows their Fourier power spectrum. The results in these figures are consistent with our analysis statements in Lemma 4.2. See Figure 57 in Appendix G.1 for similar results when $k$ is 3 or 5.
Figure 2: 2-dimension cosine shape of the trained $W^{KQ}$ (attention weights) and their Fourier power spectrum. The one-layer transformer with attention heads $m=160$ is trained on $k=4$-sum mod-$p=31$ addition dataset. We even split the whole datasets ($p^{k}=31^{4}$ data points) into training and test datasets. Every row represents a random attention head from the transformer. The left figure shows the final trained attention weights being an apparent 2-dim cosine shape. The right figure shows their 2-dim Fourier power spectrum. The results in the figures are consistent with Figure 1. See Figure 9 and Figure 10 in Appendix G.2 for similar results when $k$ is 3 or 5.
Figure 3: All Fourier spectrum frequencies being covered and the maximum normalized power of the embeddings (hidden layer weights). The one-hidden layer network with $m=2944$ neurons is trained on $k=4$-sum mod-$p=47$ addition dataset. We denote $\widehat{u}[i]$ as the Fourier transform of $u[i]$. Let $\max_{i}|\widehat{u}[i]|^{2}/(\sum|\widehat{u}[j]|^{2})$ be the maximum normalized power. Mapping each neuron to its maximum normalized power frequency, (a) shows the final frequency distribution of the embeddings. Similar to our construction analysis in Lemma 4.3, we have an almost uniform distribution over all frequencies. (b) shows the maximum normalized power of the neural network with random initialization. (c) shows, in frequency space, the embeddings of the final trained network are one-sparse, i.e., maximum normalized power being almost 1 for all neurons. This is consistent with our maximum-margin analysis results in Lemma 4.3. See Figure 6 and Figure 8 in Appendix G.1 for similar results when $k$ is 3 or 5.
@@ 4.1 Technique Overview
In this section, we propose techniques overview of the proof for our main result. We use $\mathbf{i}$ to denote $\sqrt{-1}$. Let $f:\mathbb{Z}_{p}\rightarrow\mathbb{C}$. Then, for each frequency $j\in\mathbb{Z}_{p}$, we define $f$ discrete Fourier transform (DFT) as
$\widehat{f}(j):=\sum_{\zeta\in\mathbb{Z}_{p}}f(\zeta)\exp(-2\pi\mathbf{i}\cdot j\zeta/p).$
Let ${\cal B}$ denote the ball that $\|w\|^{2}+\|u_{1}\|^{2}+\dots+\|u_{k}\|^{2}\leq 1$. Let $*\Omega_{q}^{{}^{\prime}*}$ be defined as Definition C.6. We first show how we get the single neuron class-weighted maximum-margin solution set $\Omega^{\prime*}_{q}$.
Lemma 4.2 (Informal version of Lemma C.8).
If for any $\zeta\in\{1,\ldots,\frac{p-1}{2}\}$, there exists a scaling constant $\beta\in\mathbb{R}$,
$u_{1}(a_{1})$ $=\beta\cdot\cos(\theta_{u_{1}}^{*}+2\pi\zeta a_{1}/p)$
$u_{2}(a_{2})$ $=\beta\cdot\cos(\theta_{u_{2}}^{*}+2\pi\zeta a_{2}/p)$
$\dots$
$u_{k}(a_{k})$ $=\beta\cdot\cos(\theta_{u_{k}}^{*}+2\pi\zeta a_{k}/p)$
$w(c)$ $=\beta\cdot\cos(\theta_{w}^{*}+2\pi\zeta c/p)$
Then, we have the following
$\Omega_{q}^{\prime*}=$$\{(u_{1},\dots,u_{k},w)\},$
and
$\gamma^{*}=\frac{2(k!)}{(2k+2)^{(k+1)/2}(p-1)p^{(k-1)/2}}.$
@@
The proof establishes the maximum-margin solution’s sparsity in the Fourier domain through several key steps. Initially, by Lemma C.7, focus is directed to maximizing Eq. (C.4).
For odd $p$, Eq. (C.4) can be reformulated with magnitudes and phases of $\widehat{u}_{i}$ and $\widehat{w}$ (discrete Fourie transform of $u_{i}$ and $w$), leading to an equation involving cosine of their phase differences. Plancherel’s theorem is then employed to translate the norm constraint to the Fourier domain. This allows for the optimization of the cosine term in the sum, effectively reducing the problem to maximizing the product of magnitudes of $\widehat{u}_{i}$ and $\widehat{w}$ (Eq. (26)).
By applying the inequality of arithmetic and geometric means we have an upper bound for the optimization problem. To achieve the upper bound, equal magnitudes are required for all $\widehat{u}_{i}$ and $\widehat{w}$ at a single frequency, leading to Eq. (30). The neurons are finally expressed in the time domain, demonstrating that they assume a specific cosine form with phase offsets satisfying certain conditions. ∎
See formal proof in Appendix C.3.
@@
Lemma 4.3 (Informal version of Lemma D.3).
Let $\cos_{\zeta}(x)$ denote $\cos(2\pi\zeta x/p)$. Then, we have the maximum $L_{2,k+1}$-margin solution $\theta^{*}$ will consist of $2^{2k-1}\cdot{p-1\over 2}$ neurons $*\theta^{*}_{i}\in\Omega_{q}^{{}^{\prime}*}$ to simulate $\frac{p-1}{2}$ type of cosine computation, each cosine computation is uniquely determined a $\zeta\in\{1,\ldots,\frac{p-1}{2}\}$. In particular, for each $\zeta$ the cosine computation is $\cos_{\zeta}(a_{1}+\dots+a_{k}-c),\forall a_{1},\dots,a_{k},c\in\mathbb{Z}_{p}$.
@@
Our goal is to show that $2^{2k-1}\cdot{p-1\over 2}$ neurons $*\theta_{i}^{*}\in\Omega_{q}^{{}^{\prime}*}$ are able to simulate $\frac{p-1}{2}$ type of $\cos$ computation.
We first observe that when the $\cos$ of a sum, $\cos(a_{1}+a_{2})$, is expanded, we will remove one $\cos$ product and we will add zero or two $\sin$ products. On the other hand, expanding the sine of a sum, $\sin(a_{1}+a_{2})$, we may remove one $\sin$ product and we will add one $\sin$ product as well.
[…]
By using these two observations, we have the following expansion function of $\cos_{\zeta}(x)$, which denotes $\cos(2\pi\zeta x/p)$.
$\cos_{\zeta}(\sum_{i=1}^{k}a_{i})=$ $~{}\sum_{b\in\{0,1\}^{k}}\prod_{i=1}^{k}\cos^{1-b_{i}}(a_{i})\cdot\sin^{b_{i}}(a_{i})$
$\cdot{\bf 1}[\sum_{i=1}^{k}b_{i}\%2=0]\cdot(-1)^{{\bf 1}[\sum_{i=1}^{k}b_{i}\%4=2]}.$
Note that we have $2^{k}$ terms in the above equation. By using the following fact in Lemma D.1,
@@
each term can be constructed by $2^{k-1}$ neurons. Therefore, we need $2^{k-1}2^{k}$ total neurons. To simulate $\frac{p-1}{2}$ type of simulation, we need $2^{2k-1}\frac{p-1}{2}$ neurons.
Then, using the Lemma B.1, we construct the network $\theta^{*}$. Finally, by using the Lemma B.2 from [55], we know that it is the maximum-margin solution. ∎
See formal proof in Appendix D.3. Now, we are ready to prove our main results.
@@
By Lemma 4.2, we get $\gamma^{*}$ and the single-neuron class-weighted maximum-margin solution set $*\Omega_{q}^{{}^{\prime}*}$. By satisfying Condition 3.10, we know it is used in the maximum-margin solution.
By Lemma 4.3, we can construct the network $\theta^{*}$ that uses neurons in $*\Omega_{q}^{{}^{\prime}*}$. By Lemma B.2, we know that it is the maximum-margin solution. Finally, by Lemma E.2, we know that all frequencies are covered. ∎
See formal proof in Appendix F.2.
Figure 4: Grokking (models abruptly transition from bad generalization to perfect generalization after a large number of training steps) under learning modular addition involving $k=2,3,4,5$ inputs. We train two-layer transformers with $m=160$ attention heads on $k=2,3,4,5$-sum mod-$p=97,31,11,5$ addition dataset with $50$% of the data in the training set under AdamW [46] optimizer 1e-3 learning rate and 1e-4 weight decay. We use different $p$ to guarantee the dataset sizes are roughly equal to each other. The blue curves show training accuracy and the red ones show validation accuracy. There is a grokking phenomenon in all figures. However, as $k$ increases, the grokking phenomenon becomes weak. It implies that when the ground-truth function class becomes “complicated”, the transformers need to train more steps to overfit the training datasets and the generalization tends to be better.
@@ 5 Experiments
In Section 5.1, we conduct simulation experiments to verify our analysis for $k=3,4,5$. In Section 5.2, we show that the one-layer transformer learns 2-dimensional cosine functions in their attention weights. In Section 5.3, we show the grokking phenomenon under different $k$.
@@ 5.1 One-hidden Layer Neural Network
We conduct simulation experiments to verify our analysis. In Figure 1 and Figure 3, we use SGD to train a two-layer network with $m=2944=2^{2k-2}\cdot(p-1)$ neurons, i.e., Eq. (2), on $k=4$-sum mod-$p=47$ addition dataset, i.e., Eq. (1). Figure 1 shows that the networks trained with SGD have single-frequency hidden neurons, which support our analysis in Lemma 4.2. Furthermore, Figure 3 demonstrates that the network will learn all frequencies in the Fourier spectrum which is consistent with our analysis in Lemma 4.3. Together, they verify our main results in Theorem 4.1 and show that the network trained by SGD prefers to learn Fourier-based circuits. There are more similar results when $k$ is 3 or 5 in Appendix G.1.
@@ 5.2 One-layer Transformer
$W^{P}\begin{pmatrix}W_{1}^{V\top}E\cdot\operatorname{softmax}\left({E^{\top}W_{1}^{K}W_{1}^{Q\top}E}\right)\\ \dots\\ W_{m}^{V\top}E\cdot\operatorname{softmax}\left({E^{\top}W_{m}^{K}W_{m}^{Q\top}E}\right)\end{pmatrix},$ (10)
@@
In Figure 2 , we train a one-layer transformer with $m=160$ heads attention, i.e., Eq. (10), on $k=4$-sum mod-$p=31$ addition dataset, i.e., Eq. (1). Figure 2 shows that the SGD trained one-layer transformer learns 2-dim cosine shape attention matrices, which is similar to the one-hidden layer neural networks in Figure 1. This means that the attention layer has a learning mechanism similar to neural networks in the modular arithmetic task. It prefers to learn (2-dim) Fourier-based circuits when trained by SGD. There are more similar results when $k$ is 3 or 5 in Appendix G.2.
@@ 5.3 Grokking under Different $k$
Following the experiments’ protocol in [66], we show there is the grokking phenomenon under different $k$. We train two-layer transformers with $m=160$ attention heads on $k=2,3,4,5$-sum mod-$p=97,31,11,5$ addition dataset with 50% of the data in the training. We use different $p$ to guarantee the dataset sizes are roughly equal to each other. As $k$ increases, the grokking phenomenon becomes weak. It implies that when the ground-truth function class becomes “complicated”, the transformers need to train more steps to fit the training datasets and the generalization tends to be better. This is consistent with our analysis. In Theorem 4.1, we show that we need $2^{2k-2}\cdot(p-1)$ neurons to get maximum-margin solution, and we can check this value is $10242^{2}\cdot(97-1)=384,2^{4}\cdot(31-1)=480,2^{6}\cdot(11-1)=640,2^{8}\cdot(5-1)=$ 1024 for $k=2,3,4,5$. It means that, when $k$ becomes larger, although the whole input space size is roughly unchanged, we need more neurons to fit the optimal solution as the function class becomes more “complicated”. Thus, with increasing $k$, if we train the same size of models under these settings, the models will slightly transfer from overfitting to underfitting, so the grokking phenomenon becomes weak.

### v2

@@ 3.3 Connection between Training and the Maximum Margin Solutions
Condition 3.8 (Condition C.1 in page 8 in [64]).
We have $g^{\prime}(\theta^{*},x,y)=g(\theta^{*},x,y)$ for all $(x,y)\in\operatorname{spt}(q^{*})$. It means: $\{y^{\prime}\in\mathcal{Y}\backslash\{y\}:\tau(x,y)[y^{\prime}]>0\}\subseteq\underset{y^{\prime}\in\mathcal{Y}\backslash\{y\}}{\arg\max}f(\theta^{*},x)[y^{\prime}].$
Thus, under these conditions, we only need to focus on the class-weighted maximum-margin solutions in our following analysis.
@@ 4 Main Result
We characterize the Fourier features to perform modular addition with $k$ input in the one-hidden-layer neuron network. We show that every neuron only focus on a distinct Fourier frequency. Additionally, within the network, there is at least one neuron for each frequency. When we consider the uniform class weighting, where $\mathcal{L}_{\lambda}(\theta)$ is based on $\tau(a_{1},\dots,a_{k})[c^{\prime}]:=1/(p-1)~{}~{}\forall c^{\prime}\neq a_{1}+\dots+a_{k},$ we have the following main result:
Theorem 4.1 (Main result, informal version of Theorem H.2).
Let $f(\theta,x)$ be the one-hidden layer networks defined in Eq (2). If $m\geq 2^{2k-1}\cdot\frac{p-1}{2}$, then the max $L_{2,k+1}$-margin network satisfies:
@@
    $u_{i}(a_{i})$ $=\beta\cdot\cos(\theta_{u_{i}}^{*}+2\pi\zeta a_{i}/p),~{}~{}\forall i\in[k]\quad{\text{ and }}\quad w(c)=\beta\cdot\cos(\theta_{w}^{*}+2\pi\zeta c/p),$
@@
Theorem 4.1 tells us when the number of neurons is large enough, e.g., $m\geq 2^{2k-1}\cdot\frac{p-1}{2}$ (the lower bound of $m$ may not be the tightest in our analysis), the one hidden neural network will exactly learn all Fourier frequencies/spectrum/basis to recover the modular addition operation. More specifically, each neuron will only focus on one Fourier frequency. Thus, our analysis provides a comprehensive understanding of why neural networks trained by SGD prefer to learn Fourier-based circuits. We provide the proof sketch of Theorem 4.1 in the following section.
@@ 4.1 Technique Overview
In this section, we propose techniques overview of the proof for our main result. We use $\mathbf{i}$ to denote $\sqrt{-1}$. Let $f:\mathbb{Z}_{p}\rightarrow\mathbb{C}$. Then, for each frequency $j\in\mathbb{Z}_{p}$, we define $f$ discrete Fourier transform (DFT) as $\widehat{f}(j):=\sum_{\zeta\in\mathbb{Z}_{p}}f(\zeta)\exp(-2\pi\mathbf{i}\cdot j\zeta/p).$ Let $\Omega_{q}^{{}^{\prime}*}$ be the single neuron class-weighted maximum-margin solution set (formally defined in Definition E.6). We first show how we get $\Omega^{\prime*}_{q}$.
Lemma 4.2 (Informal version of Lemma E.8).
If for any $\zeta\in\{1,\ldots,\frac{p-1}{2}\}$, there exists a scaling constant $\beta\in\mathbb{R}$, such that $u_{i}(a_{i})=\beta\cdot\cos(\theta_{u_{i}}^{*}+2\pi\zeta a_{i}/p)$ for any $i\in[k]$ and $w(c)=\beta\cdot\cos(\theta_{w}^{*}+2\pi\zeta c/p)$, where $\theta_{u_{1}}^{*},\dots,\theta_{u_{k}}^{*},\theta_{w}^{*}\in\mathbb{R}$ are some phase offsets satisfying $\theta_{u_{1}}^{*}+\dots+\theta_{u_{k}}^{*}=\theta_{w}^{*}$. Then, we have the following
$\Omega_{q}^{\prime*}=$$\{(u_{1},\dots,u_{k},w)\},\quad\gamma^{*}=\frac{2(k!)}{(2k+2)^{(k+1)/2}(p-1)p^{(k-1)/2}}.$
@@
See formal proof in Appendix E.5. The proof establishes the maximum-margin solution’s sparsity in the Fourier domain through several key steps. Initially, by Lemma E.7, focus is directed to maximizing Eq. (E.4). For odd $p$, Eq. (E.4) can be reformulated with magnitudes and phases of $\widehat{u}_{i}$ and $\widehat{w}$ (discrete Fourie transform of $u_{i}$ and $w$), leading to an equation involving cosine of their phase differences. Plancherel’s theorem is then employed to translate the norm constraint to the Fourier domain. This allows for the optimization of the cosine term in the sum, effectively reducing the problem to maximizing the product of magnitudes of $\widehat{u}_{i}$ and $\widehat{w}$ (Eq. (20)). By applying the inequality of arithmetic and geometric means we have an upper bound for the optimization problem. To achieve the upper bound, equal magnitudes are required for all $\widehat{u}_{i}$ and $\widehat{w}$ at a single frequency, leading to Eq. (24). The neurons are finally expressed in the time domain, demonstrating that they assume a specific cosine form with phase offsets satisfying certain conditions. ∎
@@
Lemma 4.3 (Informal version of Lemma F.3).
Let $\cos_{\zeta}(x)$ denote $\cos(2\pi\zeta x/p)$. Then, we have the maximum $L_{2,k+1}$-margin solution $\theta^{*}$ will consist of $2^{2k-1}\cdot{p-1\over 2}$ neurons $\theta^{*}_{i}\in\Omega_{q}^{{}^{\prime}*}$ to simulate $\frac{p-1}{2}$ type of cosine computation, each cosine computation is uniquely determined a $\zeta\in\{1,\ldots,\frac{p-1}{2}\}$. In particular, for each $\zeta$ the cosine computation is $\cos_{\zeta}(a_{1}+\dots+a_{k}-c),\forall a_{1},\dots,a_{k},c\in\mathbb{Z}_{p}$.
@@
See formal proof in Appendix F.3. Our goal is to show that $2^{2k-1}\cdot{p-1\over 2}$ neurons $\theta_{i}^{*}\in\Omega_{q}^{{}^{\prime}*}$ are able to simulate $\frac{p-1}{2}$ type of $\cos$ computation.
We have the following expansion function of $\cos_{\zeta}(x)$, which denotes $\cos(2\pi\zeta x/p)$.
$\cos_{\zeta}(\sum_{i=1}^{k}a_{i})=$ $~{}\sum_{b\in\{0,1\}^{k}}\prod_{i=1}^{k}\cos^{1-b_{i}}(a_{i})\cdot\sin^{b_{i}}(a_{i})\cdot{\bf 1}[\sum_{i=1}^{k}b_{i}\%2=0]\cdot(-1)^{{\bf 1}[\sum_{i=1}^{k}b_{i}\%4=2]}.$
The above equation can decompose a $\cos(\sum)$ to some basic elements. Note that we have $2^{k}$ terms in the above equation. By using the following fact in Lemma F.1,
@@
where each term can be constructed by $2^{k-1}$ neurons. Therefore, we need $2^{k-1}2^{k}$ total neurons. To simulate $\frac{p-1}{2}$ type of simulation, we need $2^{2k-1}\frac{p-1}{2}$ neurons. Then, using the Lemma D.1, we construct the network $\theta^{*}$. Finally, by using the Lemma D.2 from [64], we know that it is the maximum-margin solution. ∎
Now, we are ready to prove our main results.
@@
See formal proof in Appendix H.2. By Lemma 4.2, we get $\gamma^{*}$ and the single-neuron class-weighted maximum-margin solution set $\Omega_{q}^{{}^{\prime}*}$. By satisfying Condition 3.8, we know it is used in the maximum-margin solution. By Lemma 4.3, we can construct the network $\theta^{*}$ that uses neurons in $\Omega_{q}^{{}^{\prime}*}$. By Lemma D.2, we know that it is the maximum-margin solution. Finally, by Lemma G.2, we know that all frequencies are covered. ∎
@@ 5 Experiments
In Section 5.1, we conduct simulation experiments to verify our analysis for $k=3,4,5$. In Section 5.2, we show that the one-layer transformer learns 2-dimensional cosine functions in their attention weights. In Section 5.3, we show the grokking phenomenon under different $k$. Please refer to Appendix I.1 for more details about implementation.
Figure 2: All Fourier spectrum frequencies being covered and the maximum normalized power of the embeddings (hidden layer weights). The one-hidden layer network with $m=2944$ neurons is trained on $k=4$-sum mod-$p=47$ addition dataset. We denote $\widehat{u}[i]$ as the Fourier transform of $u[i]$. Let $\max_{i}|\widehat{u}[i]|^{2}/(\sum|\widehat{u}[j]|^{2})$ be the maximum normalized power. Mapping each neuron to its maximum normalized power frequency, (a) shows the final frequency distribution of the embeddings. Similar to our construction analysis in Lemma 4.3, we have an almost uniform distribution over all frequencies. (b) shows the maximum normalized power of the neural network with random initialization. (c) shows, in frequency space, the embeddings of the final trained network are one-sparse, i.e., maximum normalized power being almost 1 for all neurons. This is consistent with our max-margin analysis results in Lemma 4.3. See Figure 6 and Figure 8 in Appendix I.2 for similar results when $k$ is 3 or 5.
@@ 5.1 One-hidden Layer Neural Network
We conduct simulation experiments to verify our analysis. In Figure 1 and Figure 2, we use SGD to train a two-layer network with $m=2944=2^{2k-2}\cdot(p-1)$ neurons, i.e., Eq. (2), on $k=4$-sum mod-$p=47$ addition dataset, i.e., Eq. (1). Figure 1 shows that the networks trained with SGD have single-frequency hidden neurons, which support our analysis in Lemma 4.2. Furthermore, Figure 2 demonstrates that the network will learn all frequencies in the Fourier spectrum which is consistent with our analysis in Lemma 4.3. Together, they verify our main results in Theorem 4.1 and show that the network trained by SGD prefers to learn Fourier-based circuits. There are more similar results when $k$ is 3 or 5 in Appendix I.2.
@@ 5.2 One-layer Transformer
$W^{P}\begin{pmatrix}W_{1}^{V\top}E\cdot\operatorname{softmax}\left({E^{\top}W_{1}^{K}W_{1}^{Q\top}E}\right)\\ \dots\\ W_{m}^{V\top}E\cdot\operatorname{softmax}\left({E^{\top}W_{m}^{K}W_{m}^{Q\top}E}\right)\end{pmatrix},$
@@
In Figure 3 , we train a one-layer transformer with $m=160$ heads attention, i.e., above equation, on $k=4$-sum mod-$p=31$ addition dataset, i.e., Eq. (1). Figure 3 shows that the SGD trained one-layer transformer learns 2-dim cosine shape attention matrices, which is similar to the one-hidden layer neural networks in Figure 1. This means that the attention layer has a learning mechanism similar to neural networks in the modular arithmetic task. It prefers to learn (2-dim) Fourier-based circuits when trained by SGD. There are more similar results when $k$ is 3 or 5 in Appendix I.3.
Figure 3: 2-dimension cosine shape of the trained $W^{KQ}$ (attention weights) and their Fourier power spectrum. The one-layer transformer with attention heads $m=160$ is trained on $k=4$-sum mod-$p=31$ addition dataset. We even split the whole datasets ($p^{k}=31^{4}$ data points) into training and test datasets. Every row represents a random attention head from the transformer. The left figure shows the final trained attention weights being an apparent 2-dim cosine shape. The right figure shows their 2-dim Fourier power spectrum. The results in the figures are consistent with Figure 1. See Figure 9 and Figure 10 in Appendix I.3 for similar results when $k$ is 3 or 5.
@@ 5.3 Grokking under Different $k$
Following the experiments’ protocol in [77], we show there is the grokking phenomenon under different $k$. We train two-layer transformers with $m=160$ attention heads on $k=2,3,4,5$-sum mod-$p=97,31,11,5$ addition dataset with 50% of the data in the training. We use different $p$ to guarantee the dataset sizes are roughly equal to each other.
Figure 4 shows that the grokking weakens as the number of $k$ increases, which is consistent with our analysis. It implies that when the ground-truth function class becomes “complicated”, the transformers need to train more steps to fit the training datasets and the generalization tends to be better. Brilliant recent works by [55, 52] argue that, during learning, the network will be first in the lazy training/NTK regime and then transfer to the rich/feature learning regime sharply, leading to a grokking phenomenon. We use the learning steps required for regime switch as a metric of grokking strength.
“Underfitting” in NTK but “overfitting” in feature learning.
Note that NTK is a notorious overparameterized regime, which probably needs a much larger number of neurons than our max-margin convergence case, i.e., much larger than $\Omega(2^{2k})$ in Theorem 4.1. Thus, under the fixed $m$ and increasing $k$, the model may easily escape the NTK regime, or there is no longer an NTK regime. Thus, we will see a weaker grokking phenomenon as the learning steps needed to transfer from the NTK regime to the feature learning regime become fewer. With increasing $k$, the model will have an “underfitting” issue in the NTK regime, meaning the model must need feature learning to fit the task but cannot only fit the task by NTK. However, the model still has an “overfitting” in the feature learning regime.
Figure 4: Grokking (models abruptly transition from bad generalization to perfect generalization after a large number of training steps) under learning modular addition involving $k=2,3,4,5$ inputs. We train two-layer transformers with $m=160$ attention heads on $k=2,3,4,5$-sum mod-$p=97,31,11,5$ addition dataset with $50$% of the data in the training set under AdamW [54] optimizer 1e-3 learning rate and 1e-3 weight decay. We use different $p$ to guarantee the dataset sizes are roughly equal to each other. The blue curves show training accuracy and the red ones show validation accuracy. There is a grokking phenomenon in all figures. However, as $k$ increases, the grokking phenomenon becomes weak. See explanation in Section 5.3.